OpenStudy (anonymous):

For the rational expression, list the values that are excluded from the domain. 4/x^2-9x+20

OpenStudy (anonymous):

set the denominator = 0, solve, and those are out

OpenStudy (anonymous):

\[x^2-9x+20=0\] \[(x-5)(x-4_=0\] \[x=5\] or \[x=4\] and those values are excluded

OpenStudy (anonymous):

if its \[\frac{4}{x^2}-9x-20\ then\ only\ x=0\ isn't\ permissible.\]